# Science Park 904, 1098 XH, Amsterdam, The Netherlands

### v1

@@ Abstract
We consider the holographic description of the dynamical black hole entropy in $f(R)$ higher curvature gravity proposed by Hollands-Wald-Zhang. On the bulk side, we show that the coarse-grained entropy (outer entropy) of a generalized marginally trapped surface corresponds precisely to the Wald entropy associated with this surface. To get this result, we first formulate the AdS/CFT correspondence in the Einstein frame and derive the correspondence between von Neumann entropy of the Einstein frame and the $f(R)$ frame. This facilitates the derivation of the correspondence between the two outer entropies in the two frames. Furthermore, we directly derive a focusing theorem associated with generalized expansion in $f(R)$ gravity. We then formulate how to construct the stationary null hypersurface for the generalized expansion and the junction condition to glue a hypersurface in $f(R)$ gravity. Combining these results, we derive the expression for the entropy in the $f(R)$ frame and identify its holographic dual.
@@ 1 Introduction
where ${\mathbf{Q}}_{\xi}$ is the Noether charge of the horizon Killing field $\xi^{a}$, and $\mathbf{B}_{\mathcal{H}^{+}}$ is a specific (n1)(n-1)-form constructed on the future horizon $\mathcal{H}^{+}$. In Einstein gravity, this entropy takes the form $S_{\text{dyn}}=(1-v\partial_{v})S_{\text{BH}}$, [^1] This formula was also provided in Rignon-Bret:2023fjq along with an alternative formula which vanishes on any cross section of a light cone in Minkowski spacetime. where $v$ is an affine parameter along the horizon.
@@ 1.1 Assumptions, Conventions, and Definitions
In order to distinguish from the Einstein frame and for the convenience of subsequent descriptions, we use the $f(R)$ frame to denote the frame in which the gravitational action is $f(R)$. We use the superscripts $E$ and $f$ to emphasize which frame we are working in.
@@
Here $\theta_{k}$ is the ordinary expansion (details in section 5.1) Assume that $k^{a}=(\partial_{v})^{a}$ is affinely parametrized and $v$ is the affine parameter. We can rewrite it as
@@
    $\mu$ is homologous to $B$, and there exists a Cauchy slice $\Sigma_{min}[\mu]$ of $O_{W}^{f}[\mu]$ on which $\mu$ is a minimal Wald entropy surface homologous to $B$.
@@ 2.1 AdS/CFT Correspondence in the Einstein Frame
One can easily verify that $\Omega^{2}=(f^{\prime}(\Psi))^{\frac{2}{D-2}}>0$ (implying $\Omega>0$)). Furthermore, the Weyl transformation preserves the causal structure of the spacetime. We redefine the auxiliary field $\Psi$ by
@@
Suppose in the $f(R)$ frame, the on shell asymptotic AdS curvature is $R_{0}$. From (11), we can determine the vacuum solution
@@
    here we use the on shell condition $\Psi=R$.
@@
Thus, we obtain a one-to-one correspondence between on shell asymptotic AdS solutions in the $f(R)$ frame and the Einstein frame.
@@
Consider the on-shell solution $\Psi=R$, when we approach the asymptotic boundary, $\Omega$ approaches a constant, while $\nabla^{2}\Omega$ and $\nabla\Omega$ vanish asymptotically. Therefore, if the original spacetime is asymptotic AdS with curvature $R_{0}$, the spacetime in the Einstein frame is still asymptotic AdS, and the asymptotic scalar curvature in the Einstein frame
@@ 2.2.1 NCC Condition in the Einstein Frame
Since the Weyl transformation preserves the causal structure, the null vector $k_{E}^{a}$ is still null in the $f(R)$ frame. Therefore, if the $f(R)$ frame satisfies the null energy condition
@@ 2.2.2 Junction Conditions in the Einstein Frame
    There exists a choice of $k^{a}$ and $l^{a}$ defined on both side of $\sigma$ such that the following conditions are satisfied:
    $\begin{split}[\theta^{E}_{k}]&=0\\ [\theta^{E}_{l}]&=0\\ [\chi^{E}_{k}]&=-[\chi^{E}_{l}]=0.\end{split}$ (45)
@@ 3 Coarse-Grained Entropy of Marginally Trapped Surfaces in the Einstein Frame
In the Einstein frame, consider a (D2)(D-2)-dimensional surface $\sigma$. The outer entropy associated with $\sigma$ (homologous to $B$, the CFT boundary slice ) is defined by maximizing the von Neumann entropy over all possible inner wedge data $\{\alpha\}$, while keeping the outer wedge $O_{W}^{E}[\sigma]$ fixed Engelhardt:2018kcs :
@@
We now demonstrate that the outer entropy of a marginally trapped surface $\mu$ (satisfying the minimar condition [^4] Minimar condition: 1. $\mu$ is homologous to a connected component $B$ of CFT at one time. And there exists a Cauchy slice $\Sigma$ in $O^{E}_{W}[\mu]$ such that $\mu$ has minimal area on $\Sigma$. 2. There exists a choice of normalization for $l^{a}$ such that $\nabla^{E}_{k}\theta^{E}_{l}\leq 0$ on $\mu$, with equality allowed only if $\theta^{E}_{l}=0$ everywhere on $\mu$.) is proportional to its area in the Einstein frame. First, we show that the outer entropy of a marginally trapped surface $\mu$ is bounded from above by its area:
@@ 3.1 The Upper Bound of the Outer Entropy
We now show that $S_{E}^{\text{(outer)}}[\mu]$ is bounded by $\frac{\text{Area}[\mu]}{4G}$ in the Einstein frame. Assume the surface $\mu$ is marginally trapped. Let $N_{\pm k}(\mu)$ denote the null congruences in the Einstein frame generated from $\mu$ by the null geodesics $k_{E}^{a}$ and $k_{E}^{a}$, where $k_{E}^{a}$ is future-directed. The representative of $\mu$ on Cauchy slice $\Sigma$ Wall:2012uf ; Engelhardt:2018kcs ; Engelhardt:2017aux
$\overline{\mu}=N_{\pm k}\cap\Sigma,$ (49)
@@
Under the NCC, the area of $\mu$ is maximal on $N_{\pm k}$.
@@ 3.2 Construct the Geometry that Saturates the Upper Bound
We choose a gauge in null coordinates $u_{E},v_{E}$ and spatial coordinates $x^{i}$. Fixing $\mu$ at $u_{E}=0$ and $v_{E}=0$, we set $l^{a}=(\partial/\partial u_{E})^{a}$ and $k^{a}=(\partial/\partial v_{E})^{a}$. The gauge conditions are:
@@
$\chi^{E}_{i,v_{E}}$ $=-\theta_{v_{E}}\chi^{E}_{i}+(\frac{D-3}{D-2})\nabla_{i}\theta_{v_{E}}-(\nabla\cdot\varsigma_{v_{E}})_{i}+8\pi GT_{iv_{E}}.$ (57)
@@
Substituting (58) and (59) into (56), and noting that $\theta_{v_{E}}|_{v_{E}=0}=0$, implies $\theta_{v_{E}}=0$. Consequently, $N_{-k_{E}}$ is stationary, and $\mathcal{R}^{E}$ is also constant along $N_{-k_{E}}$. Substituting (60) into (57) reveals that $\chi_{i}$ is constant on $N_{-k_{E}}$. Together with (61) and (55), it shows that $\theta_{u_{E},v_{E}}|_{u_{E}=0}=\text{constant}$.
@@
On $N_{-k_{E}}$, $\theta_{u_{E},v_{E}}^{E}$ is constant on slices of constant $v$. From the minimar condition on $\mu$, it follows that $\theta_{u_{E},v_{E}}^{E}|_{v_{E}}<0$. In addition, we know that $\theta^{E}_{u_{E}}[\mu]<0$, this will guarantee that there will be a surface $X$ on $N_{-k}$ which $\theta^{E}_{u_{E}}=\theta^{E}_{v_{E}}=0$. This surface $X$ need not be a slice of constant $v$; generally, $X$ is defined by $v_{E}=h(x^{i})$. Our conditions ensure the existence of $h(x^{i})$. Then we apply the CPT reflection in the Einstein frame across $X$ that takes $v_{E}\to-v_{E}$, $u_{E}\to-u_{E}$, $x^{i}\to x^{i}$. It is easy to show that all quantities that are odd under the CPT reflection vanish on $X$ in our construction. Therefore, the junction conditions are automatically satisfied across $X$. Given the initial data constructed on the Cauchy slice $\Sigma=N_{-l}[\mu]\cup N_{-k}[\mu]\tilde{N}_{-l}[\tilde{\mu}$], the entire spacetime $(\mathcal{M}^{\{\alpha\}},g^{E})$ is well-defined.
@@
$\theta_{v_{E},v_{E}}=-\frac{1}{D-2}\theta_{v_{E}}^{2}-\varsigma_{v_{E}}^{2}-8\pi GT_{v_{E}v_{E}}.$ (73)
@@ 4.1 The Schwinger-Keldysh Contour and its Gravity Dual
There is a $\mathbb{Z}_{q}$ symmetry originating from the invariance under the exchange of replicas of $\rho_{A}$. Obviously, $\partial A$ is the fixed surface under the $\mathbb{Z}_{q}$ symmetry. We will assume that this symmetry is unbroken. And this symmetry will ensure that every copy of the Schwinger-Keldysh contour is the same. Now we need to construct the gravity dual of the Schwinger-Keldysh contour. We assume that the bulk fields satisfy both the cyclical symmetry ($\tau\sim\tau+2\pi iq$) and the $\mathbb{Z}_{q}$ replica symmetry.
From the preceding discussion, it is evident that the gravity dual cannot be constructed in a Euclidean manifold in the absence of time-reflection symmetry. Instead, we try to construct the gravity dual of the Schwinger-Keldysh contour in Lorentzian spacetime. It was shown that in the bulk there is a Cauchy slice $\tilde{\Sigma}_{t}$ in the Wheeler-DeWitt patch of $\Sigma_{t}$ as the bulk dual of $\Sigma_{t}$ ($\partial\tilde{\Sigma}_{t}=\Sigma_{t}$). And the bulk evolution only contains the past of the $\tilde{\Sigma}_{t}$, i.e., in $\tilde{J}^{-}[\tilde{\Sigma}_{t}]$. In other words the initial conditions are evolved forward from $t=$-$\infty$ up to $\tilde{\Sigma}_{t}$ and then we evolve back to construct the bulk Schwinger-Keldysh contour Dong:2016hjy . And there should exist a co-dimension 2 surface $e$ as the bulk dual of the $\partial A$, and it is natural to have $\partial e=\partial A$. These can be clearly represented by figure.
@@
Now we consider the case where $q\neq 1$. Defining the corresponding bulk geometry as $\mathcal{M}_{q}$, it is natural to consider a geometry that has the same cyclical symmetry ($\tau\sim\tau+2\pi iq$) and replica $\mathbb{Z}_{q}$ symmetry. And the surface $e_{q}$ the bulk correspondence of $\partial A$, should invariant under the cyclical symmetry ($\tau\sim\tau+2\pi iq$) and replica $\mathbb{Z}_{q}$ symmetry. We also assume that the partition function of boundary should be computed in the saddle point approximation of the bulk geometry $\mathcal{M}_{q}$ which satisfies all the symmetries and bulk EOM without any singularity
@@ 4.2 Correspondence of von Neumann Entropy
We first construct a replica geometry $(g_{q}^{E},\phi,\text{matter fields},\mathcal{M}_{q})$ in the Einstein frame, exhibiting cyclical and $\mathbb{Z}_{q}$ replica symmetries with respect to the codimension-2 surface $e_{q}$. And surface $e_{q}$ give a partition in Cauchy slice $\tilde{\Sigma}_{t}$ (is located in the Wheeler-DeWitt patch of $\Sigma_{t}$) whose boundary is $\Sigma_{t}$. This geometry satisfies the Einstein equation without any singularity. When $q=1$ this geometry will return to the geometry corresponding to boundary density matrix $\rho^{E}(t)$.
@@
    If in the Einstein frame $g_{q}^{E}$ $\phi$ and other matter fields satisfy the cyclical symmetry and replica $\mathbb{Z}_{q}$ symmetry, after we move into the $f(R)$ frame, the geometry ($g_{q}$, matter fields, $\mathcal{M}_{q}$) will also satisfy the cyclical symmetry and replica $\mathbb{Z}_{q}$ symmetry.
@@
    When $q=1$ the geometry will return to the geometry corresponding to boundary density matrix $\rho^{f}(t)$, and this geometry related to the geometry corresponding to $\rho^{E}(t)$ by the transformation between $f(R)$ and the Einstein frame.
@@
if the geometries correspondent to $\rho^{E}(t)$ and $\rho^{f}(t)$ are related by the transformation between the $f(R)$ frame and the Einstein frame. Since $\mathbb{Z}_{q}$ is satisfied in both frames we can write
@@
Now we apply our construction to a special case, when the spacetime satisfies the time reflection symmetry $t\to-t$ and consider the $t=0$ surface. Now we move into the Einstein frame, and from the above construction the time reflection symmetry remains unchanged. Since the time reflection symmetry, the extrinsic curvature of $t$ direction of $e_{q}$, $K^{E,t}$ will automatically be zero. We also require that the spacetime $\mathcal{M}_{q}$ contains no singularities. We will find that the extrinsic curvature of $x$ direction of $e_{q}$, $K^{E,x}$ will be zero Dong:2016hjy . Here, $x$ and $t$ are bulk coordinates correspondent to boundary Cartesian coordinates. The constraint is equivalent to the minimal surface condition Ryu:2006bv in the Einstein frame. And the von Neumann entropy is

### v2

@@ Abstract
We consider the holographic description of the dynamical black hole entropy in $f(R)$ higher curvature gravity which proposed by Hollands-Wald-Zhang. On the bulk side, we show that the coarse-grained entropy (outer entropy) of a generalized marginally trapped surface corresponds precisely to the Wald entropy associated with this surface. To get this result, we first formulate the AdS/CFT correspondence in the Einstein frame and derive the correspondence between the von Neumann entropy in the Einstein frame and the $f(R)$ frame. This facilitates the derivation of the correspondence between the two outer entropies in the two frames. Furthermore, we directly derive a focusing theorem associated with generalized expansion in $f(R)$ gravity. We then formulate how to construct the stationary null hypersurface for the generalized expansion and the junction condition to glue a hypersurface in $f(R)$ gravity. Combining these results, we directly derive the expression for the outer entropy in the $f(R)$ frame and identify its holographic dual.
@@ 1 Introduction
where ${\mathbf{Q}}_{\xi}$ is the Noether charge of the horizon Killing field $\xi^{a}$, and $\mathbf{B}_{\mathcal{H}^{+}}$ is a specific (n1)(n-1)-form constructed on the future horizon $\mathcal{H}^{+}$. In Einstein gravity, this entropy takes the form $S_{\text{dyn}}=(1-v\partial_{v})S_{\text{BH}}$ [^1] This formula was also provided in Rignon-Bret:2023fjq along with an alternative formula which vanishes on any cross section of a light cone in Minkowski spacetime., where $v$ is an affine parameter along the horizon.
@@ 1.1 Assumptions, Conventions, and Definitions
In order to distinguish from the Einstein frame and for the convenience of subsequent descriptions, we use the $f(R)$ frame to denote the frame in which the gravitational Lagrangian is $f(R)$. We use the superscripts $E$ and $f$ to emphasize which frame we are working in.
@@
Here $\theta_{k}$ is the ordinary expansion (details in section 5.1). Assume that $k^{a}=(\partial_{v})^{a}$ is affinely parametrized and $v$ is the affine parameter. We can rewrite it as
@@
    $\mu$ is homologous to $B$ (a entire connected component of the CFT at one time), and there exists a Cauchy slice $\Sigma_{min}[\mu]$ of the outer wedge $O_{W}^{f}[\mu]$ on which $\mu$ is a minimal Wald entropy surface homologous to $B$.
@@ 2.1 AdS/CFT Correspondence in the Einstein Frame
One can easily verify that $\Omega^{2}=(f^{\prime}(\Psi))^{\frac{2}{D-2}}>0$ (we assume $\Omega>0$). Furthermore, the Weyl transformation preserves the causal structure of the spacetime. We redefine the auxiliary field $\Psi$ by
@@
Suppose in the $f(R)$ frame, the on-shell asymptotic AdS curvature is $R_{0}$. From (11), we can determine the vacuum solution
@@
    here we use the on-shell condition $\Psi=R$.
@@
Thus, we obtain a one-to-one correspondence between on-shell asymptotic AdS solutions in the $f(R)$ frame and the Einstein frame.
@@
Consider the on-shell solution $\Psi=R$, when we approach the asymptotic boundary, $\Omega$ approaches a constant, while $\nabla^{2}\Omega$ and $\nabla\Omega$ vanish asymptotically. Therefore, if the original spacetime is asymptotic AdS with curvature $R_{0}$, the spacetime in the Einstein frame remains asymptotic AdS, and the asymptotic scalar curvature in the Einstein frame
@@ 2.2.1 NCC Condition in the Einstein Frame
Since the Weyl transformation preserves the causal structure, the null vector $k_{E}^{a}$ remains null in the $f(R)$ frame. Therefore, if the $f(R)$ frame satisfies the null energy condition, then
@@ 2.2.2 Junction Conditions in the Einstein Frame
    There exists a choice of $k^{a}_{E}$ and $l^{a}_{E}$ defined on both side of $\sigma$ such that the following conditions are satisfied:
    $\begin{split}[\theta^{E}_{k_{E}}]&=0\\ [\theta^{E}_{l_{E}}]&=0\\ [\chi^{E}_{k_{E}}]&=-[\chi^{E}_{l_{E}}]=0.\end{split}$ (45)
@@ 3 Coarse-Grained Entropy of Marginally Trapped Surfaces in the Einstein Frame
In the Einstein frame, consider a (D2)(D-2)-dimensional surface $\sigma$. The outer entropy associated with $\sigma$ (homologous to $B$, an entire connected component of the CFT) is defined by maximizing the von Neumann entropy over all possible inner wedge data $\{\alpha\}$, while keeping the outer wedge $O_{W}^{E}[\sigma]$ fixed Engelhardt:2018kcs :
@@
We now demonstrate that the outer entropy of a marginally trapped surface $\mu$ (satisfying the minimar condition in Einstein frame [^4] Minimar condition: 1. $\mu$ is homologous to a connected component $B$ of CFT at one time. And there exists a Cauchy slice $\Sigma$ in $O^{E}_{W}[\mu]$ such that $\mu$ has minimal area on $\Sigma$. 2. There exists a choice of normalization for $l^{a}_{E}$ such that $\nabla^{E}_{k_{E}}\theta^{E}_{l_{E}}\leq 0$ on $\mu$, with equality allowed only if $\theta^{E}_{l_{E}}=0$ everywhere on $\mu$.) is proportional to its area in the Einstein frame. First, we show that the outer entropy of a marginally trapped surface $\mu$ is bounded from above by its area:
@@ 3.1 The Upper Bound of the Outer Entropy
We now show that $S_{E}^{\text{(outer)}}[\mu]$ is bounded by $\frac{\text{Area}[\mu]}{4G}$ in the Einstein frame. Assume the surface $\mu$ is marginally trapped. Let $N_{\pm k_{E}}(\mu)$ denote the null congruences in the Einstein frame generated from $\mu$ by the null geodesics $k_{E}^{a}$ and $k_{E}^{a}$, where $k_{E}^{a}$ is future-directed. The representative of $\mu$ on Cauchy slice $\Sigma$ Wall:2012uf ; Engelhardt:2018kcs ; Engelhardt:2017aux
$\overline{\mu}=N_{\pm k_{E}}\cap\Sigma,$ (49)
@@
Under the NCC, the area of $\mu$ is maximal on $N_{\pm k_{E}}$.
@@ 3.2 Construct the Geometry that Saturates the Upper Bound
We choose a gauge in null coordinates $u_{E},v_{E}$ and spatial coordinates $x^{i}$. Fixing $\mu$ at $u_{E}=0$ and $v_{E}=0$, we set $l_{E}^{a}=(\partial/\partial u_{E})^{a}$ and $k_{E}^{a}=(\partial/\partial v_{E})^{a}$. The gauge conditions are:
@@
$\chi^{E}_{i,v_{E}}$ $=-\theta_{v_{E}}\chi^{E}_{i}+(\frac{D-3}{D-2})\nabla_{i}\theta_{v_{E}}-(\nabla\cdot\varsigma_{v_{E}})_{i}+8\pi GT_{iv_{E}},$ (57)
we will use $\theta_{u_{E}}$ and $\theta_{v_{E}}$ to emphasize our choice of gauge.
@@
Substituting (58) and (59) into (56), and noting that $\theta_{v_{E}}|_{v_{E}=0}=0$, implies $\theta_{v_{E}}=0$. Consequently, $N_{-k_{E}}$ is stationary, and $\mathcal{R}^{E}$ is also constant along $N_{-k_{E}}$. Substituting (60) into (57) reveals that $\chi^{E}_{i}$ is constant on $N_{-k_{E}}$. Together with (61) and (55), it shows that $\theta_{u_{E},v_{E}}|_{u_{E}=0}=\text{constant}$.
@@
On $N_{-k_{E}}$, $\theta_{u_{E},v_{E}}^{E}$ is constant on slices of constant $v$. From the minimar condition on $\mu$, it follows that $\theta_{u_{E},v_{E}}^{E}|_{v_{E}}<0$. In addition, we know that $\theta^{E}_{u_{E}}[\mu]<0$, this will guarantee that there will be a surface $X$ on $N_{-k}$ which $\theta^{E}_{u_{E}}=\theta^{E}_{v_{E}}=0$. This surface $X$ need not be a slice of constant $v_{E}$; generally, $X$ is defined by $v_{E}=h(x^{i})$. Our conditions ensure the existence of $h(x^{i})$. Then we apply the CPT reflection in the Einstein frame across $X$ that takes $v_{E}\to-v_{E}$, $u_{E}\to-u_{E}$, $x^{i}\to x^{i}$. It is easy to show that all quantities that are odd under the CPT reflection vanish on $X$ in our construction. Therefore, the junction conditions are automatically satisfied across $X$. Given the initial data constructed on the Cauchy slice $\Sigma=N_{-l}[\mu]\cup N_{-k}[\mu]\tilde{N}_{-l}[\tilde{\mu}$], the entire spacetime $(\mathcal{M}^{\{\alpha\}},g^{E})$ is well-defined.
@@
$\theta_{v_{E},v_{E}}=-\frac{1}{D-2}\theta_{v_{E}}^{2}-\varsigma_{v_{E}}^{2}-8\pi GT_{v_{E}v_{E}}\leq 0.$ (73)
@@ 4.1 The Schwinger-Keldysh Contour and its Gravity Dual
There is a $\mathbb{Z}_{q}$ symmetry originating from the invariance under the exchange of replicas of $\rho_{A}$. Obviously, $\partial A$ is the fixed surface under the $\mathbb{Z}_{q}$ symmetry. We will assume that this symmetry is unbroken. And this symmetry will ensure that every copy of the Schwinger-Keldysh contour is the same. Now we need to construct the gravity dual of the Schwinger-Keldysh contour. We assume that the bulk fields satisfy both the gluing condition ($\tau\sim\tau+2\pi iq$) and the $\mathbb{Z}_{q}$ replica symmetry.
From the preceding discussion, it is evident that the gravity dual cannot be constructed in a Euclidean manifold in the absence of time-reflection symmetry. Instead, we try to construct the gravity dual of the Schwinger-Keldysh contour in Lorentzian spacetime. It was shown that in the bulk there is a Cauchy slice $\tilde{\Sigma}_{t}$ in the Wheeler-DeWitt patch of $\Sigma_{t}$ as the bulk dual of $\Sigma_{t}$ ($\partial\tilde{\Sigma}_{t}=\Sigma_{t}$). And the bulk evolution only contains the past of the $\tilde{\Sigma}_{t}$, i.e., in $\tilde{J}^{-}[\tilde{\Sigma}_{t}]$. In other words the initial conditions are evolved forward from $t=$-$\infty$ up to $\tilde{\Sigma}_{t}$ and then we evolve back to construct the bulk Schwinger-Keldysh contour Dong:2016hjy . And there should exist a co-dimension 2 surface $e$ as the bulk dual of the $\partial A$, and it is natural to have $\partial e=\partial A$. These can be clearly represented by figure 2.
@@
Now we consider the case where $q\neq 1$. Defining the corresponding bulk geometry as $\mathcal{M}_{q}$, it is natural to consider a geometry that has the same gluing condition ($\tau\sim\tau+2\pi iq$) and replica $\mathbb{Z}_{q}$ symmetry. And the surface $e_{q}$ the bulk correspondence of $\partial A$, should invariant under replica $\mathbb{Z}_{q}$ symmetry. We also assume that the partition function of boundary should be computed in the saddle point approximation of the bulk geometry $\mathcal{M}_{q}$ which satisfies all the symmetries and bulk EOM without any singularity
@@ 4.2 Correspondence of von Neumann Entropy
We first construct a replica geometry $(g_{q}^{E},\phi,\text{matter fields},\mathcal{M}_{q})$ in the Einstein frame, exhibiting gluing condition and $\mathbb{Z}_{q}$ replica symmetry with respect to the codimension-2 surface $e_{q}$. And surface $e_{q}$ give a partition in Cauchy slice $\tilde{\Sigma}_{t}$ (is located in the Wheeler-DeWitt patch of $\Sigma_{t}$) whose boundary is $\Sigma_{t}$. This geometry satisfies the Einstein equation without any singularity. When $q=1$ this geometry will return to the geometry corresponding to boundary density matrix $\rho^{E}(t)$.
@@
    If in the Einstein frame $g_{q}^{E}$ $\phi$ and other matter fields satisfy the gluing condition and replica $\mathbb{Z}_{q}$ symmetry, after we move into the $f(R)$ frame, the geometry ($g_{q}$, matter fields, $\mathcal{M}_{q}$) will also satisfy the gluing condition and replica $\mathbb{Z}_{q}$ symmetry.
@@
    When $q=1$ the geometry will return to the geometry corresponding to boundary density matrix $\rho^{f}(t)$, and this geometry related to the geometry corresponding to $\rho^{E}(t)$ by the Weyl transformation between $f(R)$ and the Einstein frame.
@@
if the geometries correspondent to $\rho^{E}(t)$ and $\rho^{f}(t)$ are related by the Weyl transformation between the $f(R)$ frame and the Einstein frame. Since $\mathbb{Z}_{q}$ is satisfied in both frames we can write
@@
Now we apply our construction to a special case, when the spacetime satisfies the time reflection symmetry $t\to-t$ and consider the $t=0$ surface. Now we move into the Einstein frame, and from the above construction the time reflection symmetry remains unchanged. Since the time reflection symmetry, the extrinsic curvature of $t$ direction of $e_{q}$, $K^{E,t}$ will automatically be zero. We also require that the spacetime $\mathcal{M}_{q}$ contains no singularities. We will find that the extrinsic curvature of $x$ direction of $e_{q}$, $K^{E,x}$ will be zero Dong:2016hjy . Here, $x$ and $t$ are bulk coordinates correspondent to boundary Cartesian coordinates. The constraint is equivalent to the minimal surface condition Ryu:2006bv in the Einstein frame and the minimal Wald entropy in the $f(R)$ frame. And the von Neumann entropy is